solve for x:
59.2=42.5(1.64)^x

Found 2 solutions by Alan3354, eggsarecool:
Answer by Alan3354(6097)

59.2=42.5(1.64)^x
1.64^x = 592/425
x*log(1.64) = log(592/425)
x = log(592/425)/log(1.64)
x =~ 0.66994135

Answer by eggsarecool(32)

solve for x:

So first we divide both sides by 42.5

Now we will take ln of both sides. This helps because there is a property of logarithms that says the following In english it basically allows you to take the exponent outside the logarithm.
So applying this to our equation gives us.

Now I will move the x outside.

Now all that is left is to divide both sides by

Using a calculator you can get a decimal value of.